OpenStudy (lonnie2018):

Determine whether the lines connecting each pair of points are parallel or perpendicular. Line 1: (-4, 3) and (5, 0) Line 2: (1, -2) and (3, 4) Its supposed to be 1/3 for both but i got 1/2 for the second line.

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (usukidoll):

perpendicular lines <--- there's an intersection parallel lines <---- no intersection

OpenStudy (lonnie2018):

ii'm not doing it by graphing. its an equation.

OpenStudy (mathstudent55):

Find the slopes again. You have them both wrong.

OpenStudy (mathstudent55):

Also, what you say is supposed to be, isn't correct.

OpenStudy (mathstudent55):

Line 1: (-4, 3) and (5, 0) \(m_1 = \dfrac{0 - 3}{5 - (-4)} \) Line 2: (1, -2) and (3, 4) \(m_2 = \dfrac{4 - (-2)}{3 - 1} \)

OpenStudy (mathstudent55):

@Lonnie2018 Use my last answer and find the two slopes.

OpenStudy (usukidoll):

also parallel lines have the same slope perpendicular lines have different slopes and we use this formula \[m_1 \times m_2=-1\]